OpenStudy (kabase1234):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (sshayer):

find common difference \[a _{n}=a _{1}+(n-1)d\] \[a _{21}=a _{1}+(21-1)d,\] \[a _{1}+20d=-164\] similarly \[a _{1}+18d=-58\] subtract and find d then a1 and finally write an=?

OpenStudy (kabase1234):

An=896-53(n-1) or is it (n+1)

OpenStudy (calculusxy):

\(a_{19} = a_1 + ((19-1) \times -53)\) \(a_{19} = a_1 + (18 \times -53)\) \(-58 = a_1 -954\) \(896 = a_1\) \(a_n = 896 + (n-1)*-53\) \(a_n = 896 - 53n + 53\) \(\bf{a_n = 949 + 53n}\)

OpenStudy (kabase1234):

That can't be right

OpenStudy (mayankdevnani):

why not @Kabase1234 ?

OpenStudy (kabase1234):

My answer choices are either an=896-53(n-1) or An=896-53(n+1)

OpenStudy (calculusxy):

I meant \(949 - 53n = a_n\)

OpenStudy (calculusxy):

I simply simplified it.

OpenStudy (mayankdevnani):

this would be your answer :- an=896-53(n-1)

OpenStudy (kabase1234):

That's what I thought

OpenStudy (calculusxy):

I simplified it from: \(a_n = 896 + (n-1) * -53\) which is the same thing as \(a_n = 895-53(n-1)\)

OpenStudy (calculusxy):

\(a_n = 896 - 53(n-1)\)

OpenStudy (kabase1234):

Thank you so much! Can you help with my last one??

OpenStudy (kabase1234):

Find an equation for the nth term of the arithmetic sequence. -3, -5, -7, -9, ...

OpenStudy (kabase1234):

The second one not the first

OpenStudy (welshfella):

common difference d is -5-(-3) a1 = -3 plug these into the formula an = a1 + (n - 1)d

OpenStudy (calculusxy):

Use the main formula: \(a_n = a_1 + (n-1)d\) \(a_1\) = -3 \(d\) = -2 Substitute it back in: \(a_n = -3 + (n-1)*-2\) \(a_ n = -3 -2n + 2\) \(\bf{a_n = -1 - 2n}\)

OpenStudy (kabase1234):

-3+-2(n) or is it (n-1)

OpenStudy (kabase1234):

@calculusxy

OpenStudy (welshfella):

calculusxy has given you the correct method and answer. -3 + (n - 1)*-2 = -3 +n*-2 -1*-2 = -3 - 2n + 2 = -1 - 2n

OpenStudy (kabase1234):

That can't be the answer though

OpenStudy (kabase1234):

Choices: an = -3 - 2 an = -3 + -2(n) an = -3 + -2(n + 1) an = -3 + -2(n - 1)

OpenStudy (welshfella):

ah ok what calculusxy has done is the simplify the formula an = a1 + (n - 1)d a1 = -3 and d = -2 an = -3 + (n - 1)-2 which can be written as an = -3 + -2(n - 1)

OpenStudy (welshfella):

so that is choice d